# Totally geodesic submanifolds of the homogeneous nearly Kähler 6-manifolds and their $\mathsf{G}_{2}$-cones

### v2

@@ 1. Introduction
Although investigations about nearly Kähler manifolds began in the 1950s, significant progress has been made in recent decades. In 2005, Butruille [12] classified simply connected, homogeneous strictly nearly Kähler manifolds of dimension six. These are:
@@
More recently, in 2017, Foscolo and Haskins [19] produced the first inhomogeneous nearly Kähler structures on $\mathsf{S}^{6}$ and $\mathsf{S}^{3}\times\mathsf{S}^{3}$. These structures are of cohomogeneity one and are $\mathsf{SU}(2)\times\mathsf{SU}(2)$-invariant. The aforementioned nearly Kähler manifolds constitute all the known simply connected strictly nearly Kähler $6$-manifolds so far.
@@
The theory of totally geodesic submanifolds in homogeneous spaces is especially fruitful due to the large isometry groups of these spaces. Symmetric spaces are a particular type of homogeneous spaces with an isometric geodesic reflection at each point. Totally geodesic submanifolds of symmetric spaces have been extensively studied, see e.g. [7, 14, 15, 37]. In the setting of symmetric spaces, totally geodesic submanifolds are orbits of the action of some Lie group acting isometrically, and can be characterized algebraically by means of the so-called Lie triple systems. That being said, the classification of totally geodesic submanifolds in classical symmetric spaces of rank greater than two remains widely open. In contrast, in a non-symmetric Riemannian homogeneous space we lack a nice algebraic characterization of totally geodesic submanifolds. Furthermore, although these submanifolds are intrinsically homogeneous, they are not necessarily orbits of Lie groups acting isometrically, see [34, 48].
@@
In this article, we introduce the class of $D$-invariant totally geodesic submanifolds of a reductive homogeneous space, where $D$ denotes the difference between the Levi-Civita connection $\nabla$ and the canonical connection $\nabla^{c}$, see Subsection 2.1. In the setting of naturally reductive homogeneous spaces, which includes symmetric spaces, $D$-invariant totally geodesic submanifolds are orbits of Lie groups acting isometrically, and they admit a nice algebraic description similar to that of Lie triple systems in symmetric spaces. Furthermore, totally geodesic submanifolds of symmetric spaces are trivially $D$-invariant. Thus, the class of $D$-invariant totally geodesic submanifolds seems to be a natural generalization of totally geodesic submanifolds of symmetric spaces. As we will see, not all totally geodesic submanifolds in a naturally reductive homogeneous space are $D$-invariant. It will follow from our classification that all maximal totally geodesic submanifolds, i.e. those not contained in any other proper totally geodesic submanifold, in the homogeneous nearly Kähler $6$-manifolds are $D$-invariant.
As we have seen, there are exactly four examples of simply connected homogeneous nearly Kähler $6$-manifolds. The sphere $\mathsf{S}^{6}=\mathsf{G}_{2}/\mathsf{SU}(3)$, whose nearly Kähler structure is induced by octonionic multiplication, was the first one that appeared. Since $\mathsf{S}^{6}=\mathsf{G}_{2}/\mathsf{SU}(3)$ is isotropy irreducible, it carries the round metric. Thus, its totally geodesic submanifolds are open parts of intersections of vector subspaces of $\mathbb{R}^{7}$ passing through the center of the unit sphere $\mathsf{S}^{6}\subseteq\mathbb{R}^{7}$. Each of the remaining three examples appears as the total space of a homogeneous fibration
$F=\mathsf{K}/\mathsf{H}\rightarrow M=\mathsf{G}/\mathsf{H}\rightarrow B=\mathsf{G}/\mathsf{K}$
@@
induced by a triple of compact Lie groups $\mathsf{H}\leq\mathsf{K}\leq\mathsf{G}$. In the cases of $\mathbb{C}\mathsf{P}^{3}$ and $\mathsf{F}(\mathbb{C}^{3})$, these fibrations are also examples of twistor fibrations. A twistor fibration of an oriented Riemannian $4$-manifold $N$ is a fiber bundle $\pi\colon M\rightarrow N$, where each fiber over $p\in N$ is equal to the set of complex structures of $T_{p}N$ which preserve the orientation and the Riemannian metric of $N$. It turns out that $\pi$ is an $\mathsf{S}^{2}$-bundle over $N$, and the twistor space $M$ admits two different natural almost Hermitian structures. One of them is called the Atiyah-Hitchin-Singer structure and it is integrable if and only if the $4$-manifold $N$ is self-dual, see [4]. The other one is the Eells-Salamon structure, which can be obtained from the Atiyah-Hitchin-Singer structure by changing the sign only in the fibers. The nearly Kähler structures of $\mathbb{C}\mathsf{P}^{3}$ and $\mathsf{F}(\mathbb{C}^{3})$ that we are considering are precisely the Eells-Salamon structures that we get when $N$ is the round sphere $\mathsf{S}^{4}$ and the complex projective plane $\mathbb{C}\mathsf{P}^{2}$, respectively. Moreover, the celebrated Eells-Salamon correspondence, see [18], states a one-to-one correspondence between (branched) minimal surfaces in $N$, and non-vertical $J$-holomorphic curves in $M$, i.e. $J$-invariant immersions of $2$-manifolds in $M$.
@@
When studying the totally geodesic submanifolds of the total space of a Riemannian submersion $F\rightarrow M\rightarrow B$, it is also relevant to consider their behavior with respect to the underlying Riemannian submersion. Following [48], we say that a totally geodesic submanifold $\Sigma$ of the total space of a Riemannian submersion $M$ is well-positioned if
@@
Let $\Sigma$ be a complete submanifold of the homogeneous nearly Kähler manifold $\mathbb{C}\mathsf{P}^{3}=\mathsf{Sp}(2)/\mathsf{U}(1)\times\mathsf{Sp}(1)$ of dimension $d\geq 2$. Then, $\Sigma$ is totally geodesic if and only if it is congruent to a totally geodesic submanifold listed in Table 1.
@@
As far as we know, the totally geodesic $\mathsf{S}^{2}(\sqrt{5})$ has not appeared previously in the literature. This is an orbit of the group $\mathsf{SU}(2)_{\Lambda_{3}}$, which is the maximal connected subgroup of $\mathsf{Sp}(2)$ induced by the $4$-dimensional complex irreducible representation of $\mathsf{SU}(2)$ (note that this representation is of symplectic type). All the examples in this theorem are maximal. The non-vertical totally geodesic $J$-holomorphic curves are $\mathsf{S}^{2}(1)$ and $\mathsf{S}^{2}(\sqrt{5})$. Their associated minimal surfaces in $\mathsf{S}^{4}$ under the Eells-Salamon correspondence are a totally geodesic $2$-sphere in $\mathsf{S}^{4}$, and the Veronese embedding of the projective plane in $\mathsf{S}^{4}$, respectively.
@@
Let $\Sigma$ be a complete submanifold of the homogeneous nearly Kähler manifold $\mathsf{F}(\mathbb{C}^{3})=\mathsf{SU}(3)/\mathsf{T}^{2}$ of dimension $d\geq 2$. Then, $\Sigma$ is totally geodesic if and only if it is congruent to a totally geodesic submanifold listed in Table 2.
@@
To the best of our knowledge, $\mathbb{R}\mathsf{P}^{2}(2\sqrt{2})$ is the first example in the literature of a totally geodesic immersed submanifold of dimension $d\geq 2$ with self-intersections in a simply connected homogeneous space. Moreover, this is the only non-maximal example, it is not $D$-invariant, and not extrinsically homogeneous, i.e. an orbit of a subgroup of the isometry group of the ambient space. The non-vertical totally geodesic $J$-holomorphic curves are $\mathsf{T}^{2}_{\Lambda}$ and $\mathsf{S}^{2}(\sqrt{2})$. Their associated minimal surfaces in $\mathbb{C}\mathsf{P}^{2}$ under the Eells-Salamon correspondence are the Clifford torus in $\mathbb{C}\mathsf{P}^{2}$, that is, $\{[z_{0}:z_{1}:z_{2}]\in\mathbb{C}\mathsf{P}^{2}:\lvert z_{0}\rvert=\lvert z_{1}\rvert=\lvert z_{2}\rvert\}$, and a totally geodesic $\mathbb{R}\mathsf{P}^{2}$ in $\mathbb{C}\mathsf{P}^{2}$, respectively.
@@
Let $\Sigma$ be a complete submanifold of the homogeneous nearly Kähler manifold $\mathsf{S}^{3}\times\mathsf{S}^{3}=\mathsf{SU}(2)^{3}/\Delta\mathsf{SU}(2)$ of dimension $d\geq 2$. Then, $\Sigma$ is totally geodesic if and only if it is congruent to a totally geodesic submanifold listed in Table 3.
@@
A way to measure how a submanifold fails to be complex is by using the notion of Kähler angle, see e.g. [9]. We say that a submanifold $\Sigma$ of an almost Hermitian manifold $M$ has constant Kähler angle $\Phi(\Sigma)=\varphi\in[0,\pi/2]$ if
where $\mathrm{proj}_{T_{p}\Sigma}$ denotes the orthogonal projection onto $T_{p}\Sigma$. The submanifolds satisfying $\Phi(\Sigma)=0$ or $\Phi(\Sigma)=\pi/2$ are exactly those submanifolds which are almost complex or totally real, respectively. An interesting question is to determine the possible constant Kähler angles of the totally geodesic submanifolds of an almost Hermitian manifold. Of course, this question is only interesting for spaces with non-constant curvature, since in $\mathbb{C}^{n}$ every number in $[0,\pi/2]$ can be realized as the constant Kähler angle of a totally geodesic submanifold. In the setting of Hermitian symmetric spaces, not all totally geodesic submanifolds $\Sigma$ satisfy $\Phi(\Sigma)\in\{0,\pi/2\}$. For instance, Klein realized in [36] that there is a totally geodesic $2$-sphere in the Hermitian symmetric space $\mathsf{G}^{+}_{2}(\mathbb{R}^{5})=\mathsf{SO}(5)/(\mathsf{SO}(3)\times\mathsf{SO}(2))$ with constant Kähler angle $\arccos(1/5)$. Even more, the second author of this article proved in [50] that every rational number between $[0,1]$ can be realized as the arccosine of the Kähler angle of a totally geodesic submanifold embedded in a Hermitian symmetric space of large enough rank. As a consequence of Theorem A, Theorem B, and Theorem C, we have:
@@
Thus, every totally geodesic submanifold $\Sigma$ of a homogeneous nearly Kähler $6$-manifold of non-constant curvature satisfies $\Phi(\Sigma)\in\{0,\pi/2\}$. This raises the question whether this also holds for (not necessarily homogeneous) irreducible strictly nearly Kähler manifolds.
In this article, we also study totally geodesic submanifolds of Riemannian cones. It can be checked that for every totally geodesic submanifold $\Sigma$ of a Riemannian manifold $M$, the cone over $\Sigma$ is a totally geodesic submanifold of the cone over $M$, see Lemma 6.6. However, there might be totally geodesic submanifolds of a Riemannian cone that do not arise as cones, see Examples 6.13 and 6.15. In Section 6, we prove a structure result for totally geodesic submanifolds in cones, see Theorem 6.9. As a consequence of this, we deduce that maximal totally geodesic submanifolds of Riemannian cones are either cones over a totally geodesic submanifold or totally geodesic hypersurfaces, see Corollary 6.10. It was observed in [5] that Riemannian cones are intimately linked to special holonomy. For instance, a special class of Sasakian manifolds is that of Sasakian-Einstein manifolds, whose investigation has led to the construction of many inhomogenous Einstein metrics on spheres, see [10]. It turns out that the holonomy of the cone over a Sasakian-Einstein manifold is contained in $\mathsf{SU}(n)$. Similarly, Riemannian cones over strictly nearly Kähler $6$-manifolds have its holonomy contained in $\mathsf{G}_{2}$. This holonomy reduction is equivalent to the existence of a parallel $\mathsf{G}_{2}$-invariant three form $\phi$, i.e. a torsion-free $\mathsf{G}_{2}$-structure. Another class of $\mathsf{G}_{2}$-structures defined on $7$-dimensional manifolds is that of nearly parallel $\mathsf{G}_{2}$-structures. A $\mathsf{G}_{2}$-structure $\phi$ is nearly parallel if it satisfies $\star d\phi=c\phi$, for $c\in\mathbb{R}\setminus\{0\}$, where $\star$ denotes the Hodge star operator. It can also be proved that cones over nearly parallel $\mathsf{G}_{2}$-manifolds have its holonomy contained in $\mathsf{Spin}(7)$. Indeed, the first examples of manifolds with exceptional holonomy $\mathsf{G}_{2}$ and $\mathsf{Spin}(7)$ were constructed in [11], and they are cones over the homogeneous nearly Kähler $6$-manifold $\mathsf{F}(\mathbb{C}^{3})=\mathsf{SU}(3)/\mathsf{T}^{2}$, and the homogeneous nearly parallel $\mathsf{G}_{2}$-manifold $\mathsf{B}^{7}=\mathsf{SO}(5)/\mathsf{SO}(3)$. We prove that Sasakian-Einstein, strictly nearly Kähler $6$-manifolds, and nearly parallel $\mathsf{G}_{2}$-manifolds do not admit totally geodesic hypersurfaces, see Theorem 6.7.
@@
Moreover, Riemannian cones over $J$-holomorphic curves or Lagrangian submanifolds of a nearly Kähler $6$-manifold give rise to associative or coassociative manifolds of the corresponding $\mathsf{G}_{2}$-cone over $M$, respectively. By definition, associative and coassociative submanifolds are the submanifolds of a $\mathsf{G}_{2}$-manifold calibrated by $\phi$ and the Hodge dual of $\phi$, respectively; see [27] and [33, Chapters 4 and 12] for an introduction to calibrated geometry. As a consequence of Corollary D and Theorem E, one has the following:
@@
In Section 2, we summarize some standard facts about Riemannian homogeneous spaces (Subsection 2.1), and we describe the simply connected nearly Kähler homogeneous $6$-manifolds (Subsection 2.2). In Section 3, we recall and further develop the theory of totally geodesic immersions in analytic Riemannian manifolds. In particular, we prove a characterization of inextendable totally geodesic immersions in terms of their geodesics, see Proposition 3.4; and we make sense of the maximality notion for the case of totally geodesic immersions, see Proposition 3.6. In Section 4 we introduce and develop new tools for the study of totally geodesic submanifolds in naturally reductive homogeneous spaces. In particular, in Subsection 4.1, we consider the class of $D$-invariant totally geodesic submanifolds, and we characterize those both from an algebraic and geometric point of view, see Theorem 4.3. In Subsection 4.2, we derive a useful criterion for the existence of totally geodesic surfaces in naturally reductive homogeneous spaces, see Proposition 4.8. Also, in Subsection 4.3, we characterize when a totally geodesic submanifold of the total space of a homogeneous fibration is well-positioned, see Lemma 4.9. In Section 5, we exhibit and discuss the properties of the totally geodesic submanifolds of $\mathbb{C}\mathsf{P}^{3}$, the flag manifold $\mathsf{F}(\mathbb{C}^{3})$, and $\mathsf{S}^{3}\times\mathsf{S}^{3}$. In Section 6, we study totally geodesic submanifolds of Riemannian cones. In Subsection 6.1, we start by observing that totally geodesic submanifolds of a Riemannian manifold induce totally geodesic submanifolds in its Riemannian cone, see Lemma 6.6. We also obtain a non-existence result for totally geodesic hypersurfaces in three distinct types of manifolds: Sasakian-Einstein manifolds, six-dimensional strictly nearly Kähler manifolds, and nearly parallel $\mathsf{G}_{2}$-manifolds. Moreover, we prove Theorem 6.9, which gives a structure result for totally geodesic submanifolds in Riemannian cones. Finally, in Section 7 we provide the proofs for the main theorems.

### v3

@@ 1. Introduction
Although investigations about nearly Kähler manifolds began in the 1950s, significant progress has been made in recent decades. In 2005, Butruille [12] classified all simply connected, homogeneous, strictly nearly Kähler manifolds of dimension six. These are:
@@
More recently, in 2017, Foscolo and Haskins [19] produced the first inhomogeneous nearly Kähler structures on $\mathsf{S}^{6}$ and $\mathsf{S}^{3}\times\mathsf{S}^{3}$. These structures are of cohomogeneity one and are $\mathsf{SU}(2)\times\mathsf{SU}(2)$-invariant. The aforementioned examples constitute all the known simply connected strictly nearly Kähler $6$-manifolds so far.
@@
The theory of totally geodesic submanifolds in homogeneous spaces is especially fruitful due to the large isometry groups of these spaces. Symmetric spaces are a particular type of homogeneous spaces with an isometric geodesic reflection at each point. Totally geodesic submanifolds of symmetric spaces have been extensively studied, see e.g. [7, 13, 14, 36]. In this setting, it turns out that totally geodesic submanifolds arise as orbits of isometric actions and can be characterized algebraically by means of the so-called Lie triple systems. That being said, the classification of totally geodesic submanifolds in classical symmetric spaces of rank greater than two remains widely open. In contrast, in a non-symmetric Riemannian homogeneous space we lack a nice algebraic characterization of totally geodesic submanifolds. Furthermore, although these submanifolds are intrinsically homogeneous, they are not necessarily orbits of Lie groups acting isometrically, see [33, 46].
@@
In this article, we introduce the class of $D$-invariant totally geodesic submanifolds of a reductive homogeneous space, where $D$ denotes the difference between the Levi-Civita connection $\nabla$ and the canonical connection $\nabla^{c}$, see Subsection 2.1. In the setting of naturally reductive homogeneous spaces, which includes symmetric spaces, $D$-invariant totally geodesic submanifolds are orbits of Lie groups acting isometrically, and they admit a nice algebraic description similar to that of Lie triple systems in symmetric spaces. Furthermore, totally geodesic submanifolds of symmetric spaces are trivially $D$-invariant. Thus, the class of $D$-invariant totally geodesic submanifolds seems to be a natural generalization of totally geodesic submanifolds of symmetric spaces. As we will see, not all totally geodesic submanifolds in a naturally reductive homogeneous space are $D$-invariant. It will follow from our classification that all maximal totally geodesic submanifolds—that is, those that are not contained in any other proper totally geodesic submanifold—in the homogeneous nearly Kähler $6$-manifolds are $D$-invariant.
@@
As we have seen, there are exactly four examples of simply connected homogeneous strictly nearly Kähler $6$-manifolds. The sphere $\mathsf{S}^{6}=\mathsf{G}_{2}/\mathsf{SU}(3)$, whose nearly Kähler structure is induced by octonionic multiplication, was the first one to appear. Since $\mathsf{S}^{6}=\mathsf{G}_{2}/\mathsf{SU}(3)$ is isotropy irreducible, it carries a unique $\mathsf{G}_{2}$-invariant metric up to scaling, which is precisely the round one. Thus, its totally geodesic submanifolds are open parts of intersections of vector subspaces of $\mathbb{R}^{7}$ passing through the center of the unit sphere $\mathsf{S}^{6}\subseteq\mathbb{R}^{7}$. Each of the remaining examples appears as the total space of a homogeneous fibration $F=\mathsf{K}/\mathsf{H}\rightarrow M=\mathsf{G}/\mathsf{H}\rightarrow B=\mathsf{G}/\mathsf{K}$ induced by a triple of compact Lie groups $\mathsf{H}\subseteq\mathsf{K}\subseteq\mathsf{G}$. In the cases of $\mathbb{C}\mathsf{P}^{3}$ and $\mathsf{F}(\mathbb{C}^{3})$, these fibrations are also examples of twistor fibrations. A twistor fibration of an oriented Riemannian $4$-manifold $N$ is a fiber bundle $\pi\colon M\rightarrow N$, where each fiber over $p\in N$ is equal to the set of complex structures of $T_{p}N$ which preserve the orientation and the Riemannian metric of $N$. It turns out that $\pi$ is an $\mathsf{S}^{2}$-bundle over $N$, and the twistor space $M$ admits two different natural almost Hermitian structures. One of them is called the Atiyah–Hitchin–Singer structure and it is integrable if and only if the $4$-manifold $N$ is self-dual, see [4]. The other one is the Eells–Salamon structure, which can be obtained from the Atiyah–Hitchin–Singer structure by changing the sign only in the fibers. The nearly Kähler structures of $\mathbb{C}\mathsf{P}^{3}$ and $\mathsf{F}(\mathbb{C}^{3})$ that we are considering are precisely the Eells–Salamon structures that we get when $N$ is the round sphere $\mathsf{S}^{4}$ and the complex projective plane $\mathbb{C}\mathsf{P}^{2}$, respectively. Moreover, the celebrated Eells–Salamon correspondence, see [18], states a one-to-one correspondence between (branched) minimal surfaces in $N$, and non-vertical $J$-holomorphic curves in $M$, i.e. $J$-invariant immersions of $2$-manifolds in $M$.
@@
When studying totally geodesic submanifolds of the total space $M$ of a Riemannian submersion $F\rightarrow M\rightarrow B$, it is also relevant to consider their behavior with respect to the underlying Riemannian submersion. Following [46], we say that a totally geodesic submanifold $\Sigma$ of the total space of a Riemannian submersion $M$ is well-positioned if
@@
We are also interested in understanding the interaction between totally geodesic submanifolds of the nearly Kähler spaces under investigation and their ambient almost complex structure. A way to measure how a submanifold fails to be complex is by using the notion of Kähler angle, see e.g. [9]. We say that a submanifold $\Sigma$ of an almost Hermitian manifold $M$ has constant Kähler angle $\Phi(\Sigma)=\varphi\in[0,\pi/2]$ if
where $\mathrm{proj}_{T_{p}\Sigma}$ denotes the orthogonal projection onto $T_{p}\Sigma$. The submanifolds satisfying $\Phi(\Sigma)=0$ or $\Phi(\Sigma)=\pi/2$ are exactly those which are almost complex or totally real, respectively. An interesting question is to determine the possible constant Kähler angles of the totally geodesic submanifolds of an almost Hermitian manifold. Of course, this question is only interesting for spaces with non-constant curvature, since in $\mathbb{C}^{n}$ every number in $[0,\pi/2]$ can be realized as the constant Kähler angle of a totally geodesic submanifold. In the setting of Hermitian symmetric spaces, not all totally geodesic submanifolds $\Sigma$ satisfy $\Phi(\Sigma)\in\{0,\pi/2\}$. For instance, Klein realized in [35] that there is a totally geodesic $2$-sphere in the Hermitian symmetric space $\mathsf{G}^{+}_{2}(\mathbb{R}^{5})=\mathsf{SO}(5)/(\mathsf{SO}(3)\times\mathsf{SO}(2))$ with constant Kähler angle $\arccos(1/5)$. Even more, the second author of this article proved in [48] that every rational number between $[0,1]$ can be realized as the arccosine of the Kähler angle of a totally geodesic submanifold embedded in a Hermitian symmetric space of large enough rank.
@@
Let $\Sigma$ be a complete submanifold of the homogeneous nearly Kähler manifold $\mathbb{C}\mathsf{P}^{3}=\mathsf{Sp}(2)/\mathsf{U}(1)\times\mathsf{Sp}(1)$ of dimension $d\geq 2$. Then, $\Sigma$ is totally geodesic if and only if it is congruent to one of the submanifolds listed in Table 1.
@@
As far as we know, the totally geodesic $\mathsf{S}^{2}(\sqrt{5})$ has not appeared previously in the literature. This is an orbit of the group $\mathsf{SU}(2)_{\Lambda_{3}}$, which is the maximal connected subgroup of $\mathsf{Sp}(2)$ induced by the $4$-dimensional complex irreducible representation of $\mathsf{SU}(2)$ (note that this representation is of symplectic type). All the examples in this theorem are maximal. The non-vertical totally geodesic $J$-holomorphic curves are $\mathsf{S}^{2}(1)$ and $\mathsf{S}^{2}(\sqrt{5})$. Their associated minimal surfaces in $\mathsf{S}^{4}$ under the Eells–Salamon correspondence are a totally geodesic $2$-sphere in $\mathsf{S}^{4}$, and the Veronese embedding of the projective plane in $\mathsf{S}^{4}$, respectively.
@@
Let $\Sigma$ be a complete submanifold of the homogeneous nearly Kähler manifold $\mathsf{F}(\mathbb{C}^{3})=\mathsf{SU}(3)/\mathsf{T}^{2}$ of dimension $d\geq 2$. Then, $\Sigma$ is totally geodesic if and only if it is congruent to one of the submanifolds listed in Table 2.
@@
To the best of our knowledge, $\mathbb{R}\mathsf{P}^{2}(2\sqrt{2})$ is the first example in the literature of a totally geodesic immersed submanifold of dimension $d\geq 2$ with self-intersections in a simply connected homogeneous space. Moreover, this is the only non-maximal example, it is not $D$-invariant, and not extrinsically homogeneous, i.e. an orbit of a subgroup of the isometry group of the ambient space. The non-vertical totally geodesic $J$-holomorphic curves are $\mathsf{T}^{2}_{\Lambda}$ and $\mathsf{S}^{2}(\sqrt{2})$. Their associated minimal surfaces in $\mathbb{C}\mathsf{P}^{2}$ under the Eells–Salamon correspondence are the Clifford torus in $\mathbb{C}\mathsf{P}^{2}$, that is, $\{[z_{0}:z_{1}:z_{2}]\in\mathbb{C}\mathsf{P}^{2}:\lvert z_{0}\rvert=\lvert z_{1}\rvert=\lvert z_{2}\rvert\}$, and a totally geodesic $\mathbb{R}\mathsf{P}^{2}$ in $\mathbb{C}\mathsf{P}^{2}$, respectively.
@@
Let $\Sigma$ be a complete submanifold of the homogeneous nearly Kähler manifold $\mathsf{S}^{3}\times\mathsf{S}^{3}=\mathsf{SU}(2)^{3}/\Delta\mathsf{SU}(2)$ of dimension $d\geq 2$. Then, $\Sigma$ is totally geodesic if and only if it is congruent to one of the submanifolds listed in Table 3.
@@
As a consequence of Theorem A, Theorem B, and Theorem C, we have:
@@
Thus, every totally geodesic submanifold $\Sigma$ of a homogeneous strictly nearly Kähler $6$-manifold of non-constant curvature satisfies $\Phi(\Sigma)\in\{0,\pi/2\}$. This raises the question whether this also holds for (not necessarily homogeneous) irreducible strictly nearly Kähler manifolds.
In this article, we also study totally geodesic submanifolds of Riemannian cones. It can be checked that for every totally geodesic submanifold $\Sigma$ of a Riemannian manifold $M$, the cone over $\Sigma$ is a totally geodesic submanifold of the cone over $M$, see Lemma 6.6. However, there might be totally geodesic submanifolds of a Riemannian cone that do not arise as cones, see Examples 6.13 and 6.15. In Section 6, we prove a structure result for totally geodesic submanifolds in cones, see Theorem 6.9. As a consequence of this, we deduce that maximal totally geodesic submanifolds of Riemannian cones are either cones over a totally geodesic submanifold or totally geodesic hypersurfaces, see Corollary 6.10. It was observed in [5] that Riemannian cones are intimately linked to special holonomy. For instance, a special class of Sasakian manifolds is that of Sasakian–Einstein manifolds, whose investigation has led to the construction of many inhomogeneous Einstein metrics on spheres, see [10]. It turns out that the holonomy of the cone over a Sasakian–Einstein manifold is contained in $\mathsf{SU}(n)$. Similarly, Riemannian cones over strictly nearly Kähler $6$-manifolds have its holonomy contained in $\mathsf{G}_{2}$. This holonomy reduction is equivalent to the existence of a parallel $3$-form $\phi$ that is locally modeled on the associative $3$-form on $\mathbb{R}^{7}$, or alternatively, the existence of a torsion-free $\mathsf{G}_{2}$-structure. Another class of $\mathsf{G}_{2}$-structures defined on $7$-dimensional manifolds is that of nearly parallel $\mathsf{G}_{2}$-structures. A $\mathsf{G}_{2}$-structure $\phi$ is nearly parallel if it satisfies $\star d\phi=c\phi$, for $c\in\mathbb{R}\setminus\{0\}$, where $\star$ denotes the Hodge star operator. It can also be proved that cones over nearly parallel $\mathsf{G}_{2}$-manifolds have its holonomy contained in $\mathsf{Spin}(7)$. The first example of a manifold with exceptional holonomy $\mathsf{Spin}(7)$ was constructed in [11], and it is the cone over the homogeneous nearly parallel $\mathsf{G}_{2}$-manifold $\mathsf{B}^{7}=\mathsf{SO}(5)/\mathsf{SO}(3)$. We prove that Sasakian–Einstein, strictly nearly Kähler $6$-manifolds, and nearly parallel $\mathsf{G}_{2}$-manifolds do not admit totally geodesic hypersurfaces, see Theorem 6.7.
@@
Moreover, Riemannian cones over $J$-holomorphic curves or Lagrangian submanifolds of a nearly Kähler $6$-manifold give rise to associative or coassociative manifolds of the corresponding $\mathsf{G}_{2}$-cone over $M$, respectively. By definition, associative and coassociative submanifolds are the submanifolds of a $\mathsf{G}_{2}$-manifold calibrated by $\phi$ and the Hodge dual of $\phi$, respectively; see [26] and [32, Chapters 4 and 12] for an introduction to calibrated geometry. As a consequence of Corollary D and Theorem E, one has the following:
@@
In Section 2, we summarize some standard facts about Riemannian homogeneous spaces (Subsection 2.1), and we describe the simply connected nearly Kähler homogeneous $6$-manifolds (Subsection 2.2). In Section 3, we recall and further develop the theory of totally geodesic immersions in analytic Riemannian manifolds. In particular, we prove a characterization of inextendable totally geodesic immersions in terms of their geodesics, see Proposition 3.4; and we make sense of the maximality notion for the case of totally geodesic immersions, see Proposition 3.6. In Section 4 we introduce and develop new tools for the study of totally geodesic submanifolds in naturally reductive homogeneous spaces. In particular, in Subsection 4.1, we consider the class of $D$-invariant totally geodesic submanifolds, and we characterize those both from an algebraic and geometric point of view, see Theorem 4.3. In Subsection 4.2, we derive a useful criterion for the existence of totally geodesic surfaces in naturally reductive homogeneous spaces, see Proposition 4.8. Also, in Subsection 4.3, we characterize when a totally geodesic submanifold of the total space of a homogeneous fibration is well-positioned, see Lemma 4.9. In Section 5, we exhibit and discuss the properties of the totally geodesic submanifolds of $\mathbb{C}\mathsf{P}^{3}$, the flag manifold $\mathsf{F}(\mathbb{C}^{3})$, and $\mathsf{S}^{3}\times\mathsf{S}^{3}$. In Section 6, we study totally geodesic submanifolds of Riemannian cones. In Subsection 6.1, we start by observing that totally geodesic submanifolds of a Riemannian manifold induce totally geodesic submanifolds in its Riemannian cone, see Lemma 6.6. We also obtain a non-existence result for totally geodesic hypersurfaces in three distinct types of manifolds: Sasakian–Einstein manifolds, six-dimensional strictly nearly Kähler manifolds, and nearly parallel $\mathsf{G}_{2}$-manifolds. Moreover, we prove Theorem 6.9, which gives a structure result for totally geodesic submanifolds in Riemannian cones. Finally, in Section 7 we provide the proofs for the main theorems.